7.51A bungee cord is 30.0 m long and. when stretched a
Answer #1

After jumping off platform:
At the point where bungee becomes its full un-stretched length,
the ground is 45.0 - 30.0 = 15.0 m away.
The falling father-in-law is at this point (assuming Free-Fall conditions) moving at:
V = √2g(30.0) = √588.6 = 24.26108 m/s and has a KE = 1/2(100.0)(24.261)² = 29430 J.

The selected bungee must therefore absorb this KE
and
the add'l GPE loss {mgx} in 41.0 - 30.0 = 11.0 m. <= x-value in mgx
29430 + (100.0)(9.81)(11.0) = SPE = 1/2k(11.0)²
29430 + 10791 = (0.5)k(121) = 60.5k
k = 40221/60.5 = 664.8 N/m
when pulled with a force = 410 N, the correct bungee would stretch z: F = kz
z = 410/k = 410/664.8 = 0.617 m = 61.7 cm